OpenStudy (anonymous):

4cosxsiny=1 what is the differentiation?

OpenStudy (anonymous):

product plus chain rule

OpenStudy (anonymous):

\[4(\cos(x)\sin(y)y'-\sin(x)\sin(y))=0\]

OpenStudy (anonymous):

solve for \(y'\)

OpenStudy (anonymous):

\[4\left[ \cos x( \cos y) y' -\sin x \left( \sin y \right)\right]=0\] \[y'=\frac{ \sin x \sin y }{ \cos x \cos y }=\tan x \tan y \]